CAT Chat Minutes

August 11, 2003

 

General Information

Operation Stats - B. Wesolowski:  Information from Run 2-03:

Delivered Beam:  1454.66 hrs.
Percentage of Scheduled time:  95.20%
Downtime during Period:  73.28 hrs.
Mean Time Between Faults (MTBF):  39.32 hrs.

Faults:

#70 08/05 16:00 To 08/06 06:23 14.38

33ID rad.mon.fault [HP]

0.16 reset - refill

#71 08/06 06:32 To 08/11 07:59 121.45

Int Dump: End of Period

0.00

Beamline Downtime Report - B. Wesolowski:

Beamline Downtime Summary, 8/1/03 00:00 to 8/7/03 23:59

Date

Time

Beamline

Description

Downtime

Group

8/3/03

08:54

 

BLEPS reset requiring assistance from beamline personnel.

1:10

Beamline

8/5/03

22:37

04-ID

CPU would not change current in AC mode. Jumper installed.

0:24

XFE

8/6/03

01:35

 

Vacuum trip at both beamlines (cause under investigation).

0:25

Beamline

8/6/03

13:57

 

Air leak to the P8 shutter. Repair scheduled for shutdown.

2:00

Beamline

8/6/03

20:45

 

BLEPS reset requiring instruction from beamline personnel.

0:15

Beamline

8/7/03

14:30

33-ID

DP trip on the P5 white beam stop. ME adjusted water flow and will change filters during shutdown.

1:00

ME

8/7/03

19:15

 

User did not have ESAF describing experiment.

1:30

Beamline

Safety - E. Chang:  There will be a machine shop course held Tuesday, August 12th at 10:00 a.m. in building 401, room A1100.

Power Outage Notification - B. Wesolowski:  Some beamline staff expressed desire to be notified of power outage via e-mail to pager. This will be done for those with ANL pagerswhether this can be done for cell phones and wide-area pagers is unknown. EFOG will be contacting those on power outage list with this regard. The following e-mail notification example will be sent when power is not available:

From: MCR_Operations<>

To: mls\@aps.anl.gov

Subject: Power Outage

APS has experienced a power outage.

Lack of information to beamlines:  Until notified by investigating personnel, the main control room cannot give an estimate of the time it will take to reinstate service. MCR does not always get "good" or timely information from investigators.

Auto- restart of Oxford cryo-pumps - B. Wesolowski:  During the 6 hour power outage in July, several beamlines commented that their cryo-pumps did not restart (even though they are on emergency power and have the Oxford restart package). Viewgraphs were presented showing the location of the restart package and suggested settings of the timer relay modules. These modules have a non-changeable battery pack with a life expectancy of 10 years, however, we have found that some units have failed. It is suggested that these modules be checked. If anyone has a question regarding the location, settings, or testing, please contact Bill Wesolowski at 9496.

Preliminary schedule of work during August-September Maintenance Period B. Wesolowski:

 

Power to LOM Main Door Card Key System - B. Wesolowski: Service request is in place to put LOM main door card key system on UPS.

LOM "Fire" Door B. Wesolowski: Originally the LOM hallway fire doors were put in since LOMs were wide open and this would isolate the pentagons. Since office spaces were built, the need for isolating no longer exists. Door stops can be used to hold "fire" doors open. Eventually a new mechanism will be installed.

Emergency Power to Experiment Hall Floor - B. Wesolowski: Phase one is well in progress.